### v1

# Lossy Compression with Data, Perception, and Classification Constraints

Abstract
Balancing diverse task objectives under limited rate is crucial for developing robust multi-task deep learning (DL) models and improving performance across various domains. In this paper, we consider the lossy compression problem with human-centric and task-oriented metrics, such as perceptual quality and classification accuracy. We investigate two ternary relationships, namely, the rate-distortion-classification (RDC) and rate-perception-classification (RPC). For both RDC and RPC functions, we derive the closed-form expressions of the optimal rate for both binary and Gaussian sources. Notably, both RDC and RPC relationships exhibit distinct characteristics compared to the previous RDP tradeoff proposed by Blau et al. Then, we conduct experiments by implementing a DL-based image compression framework, incorporating rate, distortion, perception, and classification constraints. The experimental results verify the theoretical characteristics of RDC and RPC tradeoffs, providing information-theoretical insights into the design of loss functions to balance diverse task objectives in deep learning.
I Introduction
Towards an information-theoretical explanation of deep learning (DL) algorithms, the information bottleneck (IB) principle[1] has received significant attention. The IB principle introduces an additional variable that determines relevance, aiming to find a maximally compressed mapping of the input variable that preserves relevant information. This principle is connected to a remote lossy source-coding problem with logarithmic loss distortion[2, 3]. Since its introduction in 1999, the IB principle has been adapted to a broad spectrum of application domains in machine learning[4], offering a novel paradigm for analyzing deep neural networks on their layered structure[4, 5]. Meanwhile, the idea of IB has been successfully introduced to the design of communication-efficient edge inference systems[6, 7].
While the IB principle primarily focuses on tasks of target label inference, recent research has also introduced the concept of perception quality to rate-distortion analysis, thereby enabling an information-theoretical understanding of generative tasks. In the study conducted by [8], a mathematical proxy to measure perception quality based on the divergence between the source and reconstruction distributions is employed. The authors proposed the rate-distortion-perception (RDP) function, analyzing the ternary tradeoff between rate, distortion, and perception. The follow-up work [9] derived the closed-form expression of RDP function for the Gaussian source under squared error distortion and Wasserstein-2 perception loss. Meanwhile, different coding schemes have been established[10, 11, 12], showing that the RDP function can be achieved using stochastic or deterministic codes.
The tradeoffs between different tasks have been explored from the perspective of signal degradation and restoration. These studies typically involve signals that are corrupted by extrinsic noise, and only a degraded version of the source signal is available for reconstruction or denoising. In the study of perception-distortion (DP) tradeoff [13], it is demonstrated that improvements in distortion do not necessarily align with improvements in visual quality. Furthermore, in [14], classification accuracy is considered along with perceptual quality and distortion. They considered the classification error rate as a representative of semantic quality and studied classification-distortion-perception (CDP) tradeoff. The experiments showed that better classification performance comes at the cost of higher distortion or worse perception.
In this paper, we extend the scope of the aforementioned research and obtain a deeper understanding of the interplay between rate, distortion, perception, and classification. Specifically, we address two open problems: the information rate-distortion-classification (RDC) function and the information rate-perception-classification (RPC) function. For both RDC and RPC functions, we derive the closed-form expressions for the optimal rate in both binary and Gaussian scenarios. Interestingly, we find that RDC, RPC, and the previous RDP[8] exhibit distinct characteristics in each trinary relationship. We also verify our theoretical results by implementing a DL-based image compression framework that incorporates diverse constraints. As new results in characterizing task-relevant information, RDC and RPC functions complement the IB theory and RDP tradeoff, and provide information-theoretical insights into designing loss functions to balance diverse task objectives in deep learning.
Table I provides a summary of the tradeoffs between distortion, perception, and classification in both lossy compression and signal restoration frameworks, as presented in previous literature and this paper.
TABLE I: Summary of tradeoffs between distortion, perception and classification tasks in lossy compression and signal restoration.
Reference Tradeoff Closed form
[15] RD Binary[15]
Gaussian[15]
[1] RC Binary[16]
Gaussian[17]
[8] RDP Binary[8]
Gaussian[9]
[13] PD Gaussian[13]
[14] CDP
Mixed Gaussian[14]
(Simulation only)
Ours RDC Binary
Gaussian
RPC Binary
Gaussian
Note: Here R, D, P, and C refers to rate, distortion,
perception, and classification respectively.
II Problem Formulation
Consider a source with observable data $X\sim p_{X}(x)$, which intrinsically contains several classification aspects formulated by variables $S_{1},\cdots,S_{K}\sim p_{S}(s_{1},\cdots,s_{K})$. The observation $X$ and the intrinsic variables are correlated and follow the joint probability distribution $p_{X,S}(x,s_{1},\cdots,s_{K})$ over $\mathcal{X}\times\mathcal{S}_{1}\times\cdots\times\mathcal{S}_{K}$. The intrinsic variable is not observable but can be inferred from $X$. For example, the observation $X$ could be a voice signal, and the interested classification variables may be the transcription of the speech, the identity of the speaker, or the gender of the speaker.
Figure 1: Illustration of task-oriented lossy compression framework.
As shown in the Fig. 1, the process of lossy compression consists of an encoder and a decoder.
  • The encoding function $f:\mathcal{X}\mapsto\mathcal{Z}$ maps the observable source $X$ to an efficient representation $Z=f(X)$.
  • The decoding function $g:\mathcal{Z}\mapsto\mathcal{\hat{X}}$ reproduces the source data $\hat{X}$ to satisfy task-oriented demands of downstream applications.
The joint distribution of $X$ and $\hat{X}$ follows $p_{X,\hat{X}}(x,\hat{x})=p_{\hat{X}|X}(\hat{x}|x)p_{X}(x)$. Meanwhile, $(S_{1},\cdots,S_{K})$, $X$ and $\hat{X}$ follow the Markov chain $(S_{1},\cdots,S_{K})\rightarrow X\rightarrow\hat{X}$.
In task-oriented lossy compression, the destination may involve various tasks upon receiving a compressed signal. These tasks can include traditional data reconstruction, generative learning tasks, or the extraction of classification labels. As a result, the following constraints arise in order to accommodate these potential applications:
1) Reconstruction constraint: We consider the following data reconstruction constraint
$\mathbb{E}[\Delta(X,\hat{X})]\leq D,$
where $\Delta:\mathcal{X}\times\hat{\mathcal{X}}\to\mathbb{R}^{+}$ is a distortion function, such as Hamming distortion and squared-error distortion, and the expectation is taken over $p_{X,\hat{X}}(x,\hat{x})=p_{\hat{X}|X}(\hat{x}|x)p_{X}(x)$. In the subsequent discussion, the term "distortion" specifically refers to the data reconstruction distortion described here.
2) Perception constraint: The perceptual quality usually refers to the degree to which an image resembles a natural image rather than a synthetic or restored image generated by an algorithm[18]. It has been demonstrated that perceptual quality is associated with the distance between the distributions of natural images and generated images [13]. Meanwhile, the underlying principle of generative adversarial network (GAN)[19] and its variants (e.g., Wasserstein GAN[20]) also involves minimizing a divergence between two distributions. Hence, in this paper, we adopt the same perception constraint as [13, 8, 9, 14], i.e.,
$d(p_{X},p_{\hat{X}})\leq P,$
where $d(\cdot,\cdot)$ is some divergence between probability distributions, such as total-variation (TV) divergence and Kulback-Leibler (KL) divergence.
3) Classification constraint: The conditional entropy $H(X|Y)$ measures the uncertainty in $X$ given information $Y$, which also serves as a lower bound for the cross-entropy loss[21], a commonly employed metric in machine learning for optimizing classification models. Thus, we adopt the following classification constraint
$H(S_{k}|\hat{X})\leq C_{k},\qquad k\in[K],$
for some $C_{k}>0$, which means the uncertainty of classification variable $S_{k}$ given the recovered source $\hat{X}$ should not surpass the level $C_{k}$. Equivalently, the constraints can be written as $I(S_{k},\hat{X})\geq C^{{}^{\prime}}_{k},\forall k\in[K],$ which indicates that a certain amount of semantic information about the relevant variable $S_{k}$ is preserved in $\hat{X}$.
Incorporating distortion, perception, and classification constraints, we define the information rate-distortion-perception-classification (RDPC) function for a source $X\sim p_{X}(x)$ as
$R(D,P,\mathbf{C})=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (1)
s.t. $\mathbb{E}[\Delta(X,\hat{X})]\leq D,$ (1a)
$d(p_{X},p_{\hat{X}})\leq P,$ (1b)
$H(S_{k}|\hat{X})\leq C_{k},\forall k\in[K],$ (1c)
where $\mathbf{C}=(C_{1},C_{2},\cdots,C_{K})$ is the allowed uncertainty of classification variables given $\hat{X}$ in classification constraints.
Connection with previous work: The proposed $R(D,P,\mathbf{C})$ function can be reduced to a series of previous theories by relaxing one or two constraints: When the classification constraint (1c) is relaxed, $R(D,P,\mathbf{C})$ is equivalent to the RDP function in [8]. If the perception constraint is also relaxed, the RDP function is reduced to the rate-distortion function $R(D)$ [15, 22]. When there exists only a single classification variable and both reconstruction (1a) and perception constraints (1b) are inactive, the proposed $R(D,P,\mathbf{C})$ is reduced to
$\min_{p_{\hat{X}|X}:I(S,\hat{X})\geq C}I(X;\hat{X})$
By introducing the Lagrange multiplier $\beta$, the above problem is equivalent to $\min_{p_{\hat{X}|X}}\{I(X;\hat{X})-\beta I(S,\hat{X})\}$, which is the information bottleneck problem[1].
In the following sections, we will delve into a theoretical analysis of two trinary relationships: rate-distortion classification (RDC) and rate-perception-classification (RPC), which have not been investigated by previous literature. We will provide closed-form expressions for both binary and Gaussian sources, and explore distinctions between different tradeoffs.
III Rate-distortion-classification Tradeoff
In many practical scenarios, reconstruction distortion is concerned jointly with classification accuracies, such as autonomous driving and face recognition in surveillance applications[23, 24]. In this section, we characterize the inherent RDC tradeoff and derive closed-form expression of the optimal rate for binary and Gaussian sources.
When perception constraint (1b) is relaxed and consider only one classification variable, we can obtain the following information rate-distortion-classification (RDC) function:
$R(D,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (2)
s.t. $\mathbb{E}[\Delta(X,\hat{X})]\leq D,$ (2a)
$H(S|\hat{X})\leq C,$ (2b)
where $S$ is a classification variable.
III-A Binary Source
For the binary source, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
Theorem 1.
For a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$), if $C<H(S_{1})$, the problem is infeasible; otherwise, the information rate-distortion-classification function with Hamming distortion is given by
$R(D,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)\qquad\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad\text{ for }\min\{D,C_{1}\}>b,\end{cases}$
where $b=\min\{\frac{a-p_{1}}{1-2p_{1}},1-\frac{a-p_{1}}{1-2p_{1}}\}$ and $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$. Here $H^{-1}:[0,1]\rightarrow[0,\frac{1}{2}]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$.
Proof.
The main idea of this proof is to first utilize the rate-distortion theory[15] and Mrs. Gerber’s Lemma [16] to find a lower bound on the optimal rate. Then we prove that this lower bound is achievable. See the details in Appendix A. ∎
From the theorem, we can observe that when the distortion constraint $D$ is relatively large, the rate $R$ becomes a function of $C$, indicating that the classification constraint becomes the primary limiting factor. Conversely, when $D$ is relatively small and becomes the main limiting factor, the rate $R$ becomes a function of the distortion constraint $D$.
In Fig. 2(a), we visualize the result of Theorem 1 by plotting the $R(D,C)$ function of a Bernoulli source as a surface (Fig. 2(a)) as well as distortion-classification curves for different rates (Fig. 2(b)). The first observation from the figures is that $R(D,C)$ function is non-increasing and convex over $D$ and $C$. When $D$ is relatively large, the value of rate is primarily determined by the value of $C$, and vice versa. When rate bits become larger, the overall curve shifts to the lower left, which means we can achieve a better classification performance and a better distortion level simultaneously.
III-B Gaussian Source
For the scalar Gaussian source, we characterize the closed-form solution for $R(D,C)$ as the following theorem.
Theorem 2.
For a scalar Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the information rate-distortion-classification function with square error distortion is given by
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}\\ \ \ \ \ \ \ \ \ \text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))\\ \ \ \ \ \ \ \ \ \text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0\ \ \ \ \ \ \text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
where $\rho=\frac{\theta_{1}}{\sigma_{s}\sigma_{x}}$ is the correlation factor of $X$ and $S$, and $h(\cdot)$ is the differential entropy for a continuous variable.
Proof.
The converse and achievability proof is provided in Appendix B. ∎
We also visualize the result of Theorem 2 by plotting the $R(D,C)$ function of a Gaussian source as a surface as well as distortion-classification curves for different rates (Fig. 2(b)). Similar properties with binary source such as convexity and non-increasing of $R(D,C)$ on $D$ and $C$ could be observed.
Figure 2: Visualization for RDC function of (a) a Bernoulli source and (b) a Gaussian source. The left figures depict $R(D,C)$ function as a surface. The colored solid lines on the figure indicate the different rate levels. The right figures show $R(D,C)$ function along distortion-classification planes, which is the two-dimensional projection of contour lines of (a).
In addition, we investigate the activeness of distortion constraint and classification constraint given different value pairs of $(D,C)$. As illustrated in Fig. 3, $D$ and $C$ are at odds with each other in an ‘antagonistic’ way where once $D$ is relatively small ($D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{2C-h(S)}))$) and the distortion constraint is active, the classification constraint would be silent, resulting in the dominance of $D$ in determining the rate. Conversely, as $D$ increases, the classification will be the only active constraint and the rate is determined only by $C$.
Figure 3: Visualization of the regions showing activeness of the distortion and classification constraints given each pair of $(D,C)$.
The phenomenon differs from the case of RDP [8][9]. As proved in [9], in Gaussian case with Wasserstein-2 perception loss, when $\sqrt{P}\geq\sigma_{X}-\sqrt{|\sigma_{x}^{2}-D|}$, the distortion constraint becomes the only active constraint, and the rate is solely determined by $D$. However, as $P$ decreases, both $P$ and $D$ are active, leading to a joint determination of the rate by both $D$ and $P$. Note that $I(X,\hat{X})=0$ is achievable for any $P$ when $D$ is not an active constraint[8].
IV Rate-perception-classification Relationship
In practice, generative tasks and classification tasks are often jointly considered in the context of conditional generative adversarial nets (GAN)[25], with great potential in controlling the categories of generated images. In this section, we aim to investigate the inherent relationship between rate-perception-classification (RPC).
Specifically, let us consider the following information rate-perception-classification (RPC) function where distortion constraint (1a) is relaxed in RDPC function and only one classification variable is considered:
$R(P,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (3)
s.t. $d(p_{X},p_{\hat{X}})\leq P,$ (3a)
$H(S|\hat{X})\leq C,$ (3b)
where $S$ is the only classification variable.
IV-A Binary source
We now find the optimal rate $R(P,C)$ required to describe a binary source with perception quality characterized by total variation divergence.
Theorem 3.
For a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$), if $C<H(S_{1})$, the problem is infeasible; otherwise, the information rate-distortion-classification function with total variation (TV) divergence is given by
$R(P,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }H(S_{1})\leq C\leq H(S),\\ 0\qquad\text{ for }C>H(S),\end{cases}$
where $b=\min\{\frac{a-p_{1}}{1-2p_{1}},1-\frac{a-p_{1}}{1-2p_{1}}\}$ and $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$.
Proof.
The conditional entropy constraint provides a lower bound on the rate. By delicately assigning the value of conditional probability $p_{\hat{X}|X}(\hat{X}=0|X=0)$ and $p_{\hat{X}|X}(\hat{X}=0|X\!=\!1)$, the lower bound is achievable with the TV divergence $d_{TV}(p_{X},p_{\hat{X}})\!=\!0$. See the details in Appendix C. ∎
The above theorem indicates that the rate only depends on the classification constraint, and there is no tradeoff between classification and perception. The proof of Theorem 3 establishes that it is always possible to find an optimal solution that achieves a TV divergence of zero while attaining the lower bound of the rate imposed by the classification constraint.
IV-B Gaussian source
The similar conclusion could be obtained for the Gaussian case, where the optimal rate is solely determined by the level of the conditional entropy constraint, and the perception constraint is always inactive. In this subsection, we take the natural base of logarithm.
Theorem 4.
For a scalar Gaussian source $X\sim\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ and a classification variable $S\sim\mathcal{N}(\mu_{s},\sigma_{s}^{2})$ with covariance $\text{Cov}(X,S)=\theta_{1}$, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the information rate-perception-classification function with KL divergence is given by
$R(P,C)=\begin{cases}-\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ \ \ \ \ \ \ \ \ \ \text{ for }\frac{1}{2}\log(1\!-\!\rho^{2})\!+\!h(S)\leq C\leq h(S);\\ 0,\ \ \ \ \ \ \text{ for }C>h(S);\end{cases}$
where $\rho=\frac{\theta_{1}}{\sigma_{s}\sigma_{x}}$ is the correlation factor of $X$ and $S$, and $h(\cdot)$ is the differential entropy for a continuous variable.
Proof.
We will prove that the lower bound given by the conditional entropy constraint is achievable with the KL divergence always being zero. See the details in Appendix D. ∎
Remark 1.
An interesting observation is that RDC, RPC, and the previous RDP[8] exhibit distinct characteristics within each trinary relationship. Specifically, for the RDP case, the distortion constraint is always active as long as $R(D,P)>0$[8], while the activeness of perception depends on its relative value with distortion. For the RDC case, under a limited rate, distortion and classification are at odds with each other in an ‘antagonistic’ way where at most one constraint could be active at the same time. Finally, when considering the perception and classification constraints, the optimal rate solely depends on the classification and is irrelevant to the perception.
V Experimental Results
In this section, we verify our theoretical results by implementing a deep-learning (DL) based image compression framework incorporating different constraints. We illustrate the characteristics of RDC and RPC tradeoffs in the MNIST dataset, which coincide with our theoretical results.
Figure 4: Illustration of rate-distortion-perception-classification network
Similar to [9], we utilize a generative adversarial network (GAN) with stochastic encoder and decoder. As shown in Fig. 4, the system consists of an encoder $E$, a decoder $G$, a discriminator $D$, and a pre-trained classifier $C$.
The output reconstruction $\hat{x}$, as well as the original image $x$, are fed to the discriminator $G$ and compute the discriminator loss. The reconstruction $\hat{x}$ is also fed to a pre-trained classifier to obtain a one-hot vector $\hat{\mathbf{s}}$ with each entry being the probability that the image belongs to the corresponding class. The true label $s$ and the predicted $\mathbf{\hat{s}}$ are used to compute the classification accuracy.
To control the compression rate $R$, we let $R$ be the upper bound $\text{dim}\times\log_{2}(L)$, where dim is the dimension of the encoder’s output, and $L$ is the number of levels used for quantizing each entry. As discussed in [8], setting $R$ to its upper bound significantly simplifies the scheme and is found to be only slightly sub-optimal [26].
For the reconstruction loss, we use MSE $\mathbb{E}(||x-\hat{x}||^{2})$ serving as the traditional distortion loss. For the perception loss, we employ the Wasserstein-1 loss to measure the distance between distributions. For the classification loss, we use the cross-entropy loss $\text{CE}(s,\mathbf{\hat{s}})$ which is shown to be an upper bound of the conditional entropy constraint[21].
In summary, when the rate $R=\text{dim}\times\log_{2}(L)$, the overall loss function of the DL-based lossy compression framework based on the RDPC function is
$\mathcal{L}=\lambda_{d}\mathbb{E}(||x-\hat{x}||^{2})+$ $\lambda_{p}W_{1}(p_{X},p_{\hat{X}})+\lambda_{c}\text{CE}(s,\mathbf{\hat{s}}),$ (4)
where $\lambda_{d},\lambda_{p},$ and $\lambda_{c}$ are hyperparameters to control weights of distortion, perception, and cross entropy losses, respectively.
V-1 RDC tradeoff
To illustrate the relationship within RDC, we set the parameters in the loss function (4) as $\lambda_{p}=0$ and train the model with a series of $\lambda_{d}$ and $\lambda_{c}$. For the MNIST dataset, rates are controlled by $R=\text{dim}\times\log_{2}(L)$ with (dim, $L$) pairs including $(3,3)$, $(3,4)$, $(4,4)$, $(5,4)$. For points with the same color (i.e., trained with the same $R$), the results show that higher classification accuracy often requires sacrificing distortion, which coincides with our theoretical results. It is also observed that $D-C$ curves shift to the left-up when $R$ increases, achieving better distortion and higher classification accuracy simultaneously.
V-2 RPC tradeoff
Similarly, by setting the parameters in the loss function (4) as $\lambda_{d}=0$, we illustrate RPC relationship as shown in Fig. 5(b). The figure shows that points with the same color (indicating the same rate $R$) are about parallel to the perception-axis, which coincides with our theoretical result that the rate only depends on the classification constraint and there is no tradeoff with perception constraint. When $R$ increases, we can obtain higher classification accuracy but the increment above $R\geq 8$ is marginal. We can see that under each level of accuracy (or rate), the potential of zero perception loss remains unaffected.
Figure 5: Rate-distortion-classification tradeoffs in MNIST dataset. The loss is computed by setting $\lambda_{p}=0$ and varying $\lambda_{d}$ and $\lambda_{c}$.
VI Conclusion
In this paper, we considered the lossy compression problem under different constraints, including traditional distortion measure, perceptual quality characterized by divergence between distributions, and classification accuracy characterized by conditional entropy. We investigated two open problems: the information rate-distortion-classification (RDC) function and the information rate-perception-classification (RPC) function, deriving the closed-form optimal rates for both binary and Gaussian sources. Experiments on the MNIST dataset validate the theoretical results, offering insights on developing robust multi-task deep learning models and balancing diverse objectives across various domains.
Appendix A Proof of Theorem 1
Recall the following information rate-distortion classification (RDC) function:
$R(D,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (2)
s.t. $\mathbb{E}[\Delta(X,\hat{X})]\leq D,$ (2a)
$H(S|\hat{X})\leq C,$ (2b)
where $S$ is a classification variable.
Consider a Bernoulli source $X$ and a classification variable $S$ with the binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$).
From the binary symmetric joint distribution, we can obtain that the marginal distribution of $X$ is $\text{Pr}(X=1)=\frac{a-p_{1}}{1-2p_{1}}\triangleq b$. Here we always choose $b$ to be less than $\frac{1}{2}$, otherwise we simply assign $\text{Pr}(X=1)=1-b$.
Note that by data-processing inequality[15, Theorem 2.8.1] and $S-X-\hat{X}$, we have
$H(S|\hat{X})\geq H(S|X)=H(X\oplus S_{1}|X)=H(S_{1}).$
Thus, when $C<H(S_{1})$, the classification constraint $H(S|\hat{X})\leq C$ would be infeasible. We assume $C\geq H(S_{1})$ in the following discussion.
We first find a lower bound of $I(X,\hat{X})$ leveraging the constraints and then show that it is achievable. By Mrs. Gerber’s Lemma proved in [16], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, where
$C_{1}\triangleq\frac{H^{-1}(C)-p_{1}}{1-2p_{1}},$ (5)
and $H^{-1}:[0,1]\rightarrow[0,\frac{1}{2}]$ denotes the inverse function of Shannon entropy for probability less than $\frac{1}{2}$. Note that $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}\leq\frac{\frac{1}{2}-p_{1}}{1-2p_{1}}=\frac{1}{2}$. Therefore, we have
$I(X;\hat{X})=H(X)-H(X|\hat{X})\geq H(b)-H(C_{1}).$ (6)
Case 1: When $D\leq b$ and $D\leq C_{1}$, since
$\mathbb{E}[\Delta(X,\hat{X})]$ $=P(X=0,\hat{X}=1)+P(X=1,\hat{X}=0)$
$=P(X\neq\hat{X})=P(X\oplus\hat{X}=1)\leq D,$
we can obtain
$I(X;\hat{X})$ $=H(X)-H(X|\hat{X})$
$=H(b)-H(\hat{X}\oplus X|\hat{X})$
$\geq H(b)-H(\hat{X}\oplus X)$
$\geq H(b)-H(D).$ (7)
Since $D\leq C_{1}\leq\frac{1}{2}$, we have $H(D)\leq H(C_{1})$. Combining (6) and (7), we have the lower bound of rate $I(X;\hat{X})\geq H(b)-H(D)$. This lower bound can be achieved by choosing $(X,\hat{X})$ to have the joint distribution given by BSC with $X=\hat{X}\oplus S_{2}$ where $S_{2}\sim\text{Bern}(D)$, since $I(X,\hat{X})=H(X)-H(X|\hat{X})=H(b)-H(D)$. Here the classification constraint is also satisfied: By $D<C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$, we have $p_{1}*D_{0}<H^{-1}(C)$ where $p_{1}*D_{0}=p_{1}(1-D_{0})+D(1-p_{1})$, so $H(S|\hat{X})=H(p_{1}*D_{0})<C$. Thus, $R(D,C)=H(b)-H(D)$ in this case.
Case 2: When $D\geq C_{1}$ and $C_{1}\leq b$, considering (6) and (7), the lower bound is $I(X;\hat{X})\geq H(b)-H(C_{1})$. This can be achieved by choosing $(X,\hat{X})$ to have the joint distribution given by BSC with $X=\hat{X}\oplus S_{2}$ where $S_{2}\sim\text{Bern}(C_{1})$. Then $I(X,\hat{X})=H(X)-H(X|\hat{X})=H(b)-H(C_{1})$, and the distortion constraint is also satisfied, since $\mathbb{E}[\Delta(x,\hat{x})]=P(X\oplus\hat{X}=1)=C_{1}\leq D$. Thus, $R(D,C)=H(b)-H(C_{1})$.
Case 3: When $\min\{D,C_{1}\}>b$, by letting $\hat{X}=0$ with probability $1$, we have $I(X;\hat{X})=H(b)-H(\hat{X}\oplus X)=H(b)-H(b)=0$. The distortion constraint is satisfied, since $\mathbb{E}[\Delta(x,\hat{x})]=P(X\oplus\hat{X}=1)=b\leq D$. Meanwhile, since $C_{1}>b$, we have $\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}>\frac{a-p_{1}}{1-2p_{1}}$ by definition, implying $H^{-1}(C)>a$. Thus, we have $H(S|\hat{X})=H(a)<C$, which means that the classification constraint is satisfied.
In summary, when $C\leq H(S_{1})$, the problem is feasible; otherwise the rate-distortion-classification function for the binary case is
$R(D,C)=\begin{cases}H(b)-H(C_{1})\qquad\text{ for }D\geq C_{1}\text{ and }C_{1}\leq b,\\ H(b)-H(D)\qquad\text{ for }D<C_{1}\text{ and }D\leq b,\\ 0\qquad\text{ for }\min\{D,C_{1}\}>b.\end{cases}$
Appendix B Proof of Theorem 2
Consider the RDC problem as below,
$R(D,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$
s.t. $\mathbb{E}[(X-\hat{X})^{2}]\leq D,$
$H(S|\hat{X})\leq C,$
where $(X,S)$ are jointly multivariate Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [9, 17], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, thus the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[15, Chapter 8], the RDC problem can be represented by
$R(D,C)=$ $\min_{\mu_{\hat{x}},\sigma_{\hat{x},}\theta_{2}}-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ (8)
s.t. $(\mu_{x}-\mu_{\hat{x}})^{2}+\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-2\theta_{2}\leq D,$ (8a)
$-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})\geq h(S)-C.$ (8b)
The first observation is that we can assume $\mu_{\hat{x}}=\mu_{x}$ without loss of optimality, since the objective function (8) and the classification constraint (8b) are irrelevant with the value of $\mu_{\hat{x}}$, and the distortion (8a) will be further reduced by choosing $\mu_{\hat{x}}=\mu_{x}$: $(\mu_{x}-\mu_{\hat{x}})^{2}+\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-2\theta_{2}\geq\sigma_{x}^{2}+\sigma_{\hat{x}}^{2}-2\theta_{2}$.
The second observation is that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. To make $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ in (8) meaningful, we have $01-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}}>$ 0, i.e.,
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}<\sigma_{x}^{2}.$
Then, the mutual information between $S$ and $\hat{X}$ is upper bounded by
$I(S,\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})$
$\leq-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}),$
making the (8b) infeasible if $C<\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$. Thus, in the following discussion, we assume $C\geq\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$.
In the case of (8b) not active, by Shannon’s classic rate-distortion theory for a Gaussian source with squared-error distortion[15, Chapter 10], the optimal rate is $\frac{1}{2}\log\frac{\sigma_{s}^{2}}{D}$ for $D\leq\sigma_{x}^{2}$, achieved by setting $\sigma_{\hat{x}}^{2}=\sigma_{x}^{2}-D$ and $\theta_{2}=\sigma_{x}^{2}-D$. For the classification constraint (8b), it is inactive if
$I(S,\hat{X})$ $=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})$
$=-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}(\sigma_{x}^{2}-D)}{\sigma_{s}^{2}\sigma_{x}^{4}})\geq h(S)-C,$
which is equivalent to
$D\leq\sigma_{x}^{2}-\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$
We now show that when $D>\sigma_{x}^{2}-\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{2C-2h(S)})$, the distortion constraint (8a) will be inactive. The classification constraint (8b) gives us a lower bound of $\frac{\theta_{2}^{2}}{\sigma_{\hat{s}}^{2}}$:
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$ (9)
Observing that the objective function (8) is an increasing function of $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}$, and by incorporating (9), we can obtain
$I(X,\!\hat{X})\!=\!-\!\frac{1}{2}\!\log(1\!\!-\!\!\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})\!\geq\!-\!\frac{1}{2}\!\log(1\!-\!\!\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}\!(1\!-\!e^{\!-2h(\!S\!)+2C})),$
with equality holds at $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{2C-2h(S)})$. Without loss of optimality, we choose $\sigma_{\hat{x}}^{2}=\theta_{2}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$ and substitute it to the distortion function:
$\mathbb{E}[(X\!\!-\!\!\hat{X})^{2}]\!=\!\sigma_{x}^{2}\!+\!\sigma_{\hat{x}}^{2}\!-\!2\theta_{2}\!=\!\sigma_{x}^{2}\!-\!\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1\!-\!e^{\!-2h(S)+2C})\!<\!D,$
i.e., the distortion constraint are also satisfied.
When $C>h(S)$ and $D>\sigma_{x}^{2}$, we can simply choose $\sigma_{\hat{x}}=0$ (i.e., $\hat{X}$ is a constant). Then we have the rate $I(X,\hat{X})=0$, and all the constraints are satisfied:
$\mathbb{E}[(X-\hat{X})^{2}]=\sigma_{x}^{2}<D,$
$h(S|\hat{X})=h(S)<C.$
In summary, if $C<\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the problem will be infeasible; otherwise, the rate-distortion-classification function with MSE distortion is given by
$R(D,C)=\begin{cases}\frac{1}{2}\log\frac{\sigma_{x}^{2}}{D}\\ \ \ \ \ \ \ \ \ \text{ for }D\leq\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ -\frac{1}{2}\log(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C}))\\ \ \ \ \ \ \ \ \ \text{ for }D>\sigma_{x}^{2}(1-\frac{1}{\rho^{2}}(1-e^{-2h(S)+2C})),\\ 0\ \ \ \ \ \ \text{ for }C>h(S)\quad\text{and}\quad D>\sigma_{x}^{2},\end{cases}$
where $\rho=\frac{\theta_{1}}{\sigma_{s}\sigma_{x}}$ is the correlation factor of $X$ and $S$, and $h(\cdot)$ is the differential entropy for a continuous variable.
Appendix C Proof of Theorem 3
Recall the following information rate-perception classification (RPC) function:
$R(P,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$ (3)
s.t. $d(p_{X},p_{\hat{X}})\leq P,$ (3a)
$H(S|\hat{X})\leq C,$ (3b)
where $S$ is the only classification variable.
Consider a Bernoulli source $X$ and a classification variable $S$ with binary symmetric joint distribution given by $S=X\oplus S_{1}$ where $S\sim\text{Bern}(a)$ and $S_{1}\sim\text{Bern}(p_{1})$ ($a,p_{1}\leq\frac{1}{2}$). Similar to the RDC case, if $C<H(S_{1})$, by data-processing inequality[15, Theorem 2.8.1], the problem is infeasible. In the following discussion, we assume $C\geq H(S_{1})$.
To derive the optimal rate, we need to find the optimal solution $p_{a}=P(\hat{X}=0|X=0)$ and $p_{b}=P(\hat{X}=0|X=1)$ to binary RPC problem (3). From the BSC assumption, we can obtain that the marginal distribution of $X$ is $\text{Pr}(X=1)=\frac{a-p_{1}}{1-2p_{1}}\triangleq b$.
Take the perception constraint to be the total-variation (TV) divergence, i.e.,
$d_{TV}(p_{X},p_{\hat{X}})$ $=\frac{1}{2}\sum_{z\in\{0,1\}}|p_{X}(z)-p_{\hat{X}}(z)|$
$=|-bp_{b}+(1-b)(1-p_{a})|.$
Now we express $H(S|\hat{X})$ as a function of $p_{a},p_{b}$. By the definition of conditional entropy and $S-X-\hat{X}$, we have
$H(S|$ $\hat{X})\!\!=\!-(bp_{1}p_{b}\!\!+\!\!(\!1\!-\!b\!)\!(\!1\!-\!p_{1}\!)p_{a})\!\log\!\Big{(}\!\frac{bp_{1}p_{b}\!\!+\!\!(\!1\!\!-\!\!b)\!(1\!\!-\!p_{1}\!)p_{a}}{bp_{b}\!\!+\!\!(1\!\!-\!b)p_{a}}\!\Big{)}$
-- $\!(bp_{1}\!(\!1\!-\!p_{b}\!)\!\!+\!\!(1\!-\!b)(1\!-\!p_{1})(1\!-\!p_{a}\!))$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \cdot\log\Big{(}\!\frac{bp_{1}\!(1\!-\!p_{b})\!\!+\!(1\!\!-\!b)(1\!\!-\!p_{1})(\!1\!\!-\!p_{a})}{b(\!1\!-\!p_{b})\!+\!(1\!-\!b)p_{a}}\Big{)}$
-- $\!(b(1\!\!-\!p_{1})p_{b}\!+\!(1\!\!-\!b)p_{1}p_{a})\cdot\!\log\Big{(}\!\frac{b(1\!\!-\!p_{1}\!)p_{b}\!+\!(1\!-\!b)p_{1}p_{a}}{bp_{b}\!+\!(1\!-\!b)p_{a}}\!\Big{)}$
-- $(b(1\!-\!p_{1})(1\!\!-\!p_{b})\!+\!(1\!\!-\!b)p_{1}(1\!\!-\!p_{a}))$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \cdot\log\Big{(}\!\frac{b(1\!-\!p_{1})(1\!-\!p_{b})\!+\!(1\!-\!b)p_{1}(1\!-\!p_{a})}{b(1\!-\!p_{b})\!+\!(1\!-\!b)p_{a}}\!\Big{)}$
$\triangleq f(p_{a},p_{b}).$
Consider the solution
$p_{b}=\frac{(1-b)(1-p_{a})}{b},$ (10)
$p_{a}=g^{-1}(C),$ (11)
where $g^{-1}:[0,1]\rightarrow[\frac{1}{2},1]$ denote the inverse function of
$g(p_{a})\triangleq f(p_{a},\frac{(1-b)(1-p_{a})}{b}).$ (12)
First we show that for any $H(S_{1})\leq C\leq H(S)$, we can always find a solution $p_{a}\geq\frac{1}{2}$ to $g(p_{a})=C$, i.e., $g^{-1}(C)$ always exists. To simplify the notation, let
$A=(1-2p_{1})(1-b)p_{a}+p_{1}(1-b),$
$B=(2p_{1}-1)(1-b)p_{a}+p_{1}b+(1-b)(1-2p_{1}),$
$C=(2p_{1}-1)(1-b)p_{a}+(1-p_{1})(1-b),$
$D=(1-2p_{1})(1-b)p_{a}+(1-p_{1})b+(1-b)(2p_{1}-1).$
Note that $A+C=1-b$, $B+D=b$. Then we have,
$g(p_{a})$ $=-A\log\frac{A}{1-b}-B\log\frac{B}{b}-C\log\frac{C}{1-b}-D\log\frac{D}{b}$
$=(-A\log A-B\log B-C\log C-D\log D)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +(A+C)\log(1-b)+(B+D)\log b$
$=H(A,B,C,D)-H(b),$
where $H(A,B,C,D)$ is the entropy with quaternary probability distribution $\{A,B,C,D\}$, and is also a function of $p_{a}$.
By taking the derivative of $g(p_{a})$ over $p_{a}$, we have
$\frac{\text{d}g(p_{a})}{\text{d}p_{a}}$ $=-(\log A+1)\frac{\partial A}{\partial p_{a}}-(\log B+1)\frac{\partial B}{\partial p_{a}}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ -(\log C+1)\frac{\partial C}{\partial p_{a}}-(\log D+1)\frac{\partial D}{\partial p_{a}}$
$=-(1\!-\!2p_{1})(1\!-\!b)\log A\!-\!(2p_{1}\!-\!1)(1\!-\!b)\log B$
$\ \ \ \ \ \ \ \!-\!(2p_{1}\!-\!1)(1\!-\!b)\log C\!-\!(1\!-\!2p_{1})(1\!-\!b)\log D$
$=(1-2p_{1})(1-b)\log\frac{B\cdot C}{A\cdot D}.$
By solving $\frac{\text{d}g(p_{a})}{\text{d}p_{a}}=0$, i.e., $B\cdot C=A\cdot D$, we can obtain $p_{a}=1-b$. When $p_{a}<1-b$, $\frac{\text{d}g(p_{a})}{\text{d}p_{a}}>0$, $g(p_{a})$ increases; and when $1-b\leq p_{a}\leq 1$, $\frac{\text{d}g(p_{a})}{\text{d}p_{a}}\leq 0$, $g(p_{a})$ decreases. Hence, $g(p_{a})$ reaches the maximum when $p_{a}=1-b$ and
$g(1-b)=H(b*p_{1})+H(b)-H(b)=H(S).$
Meanwhile, since
$g(1)$ $=H\big{(}(1-p_{1})(1-b),p_{1}b,p_{1}(1-b),(1-p_{1})b\big{)}-H(b)$
$=H(p_{1})+H(b)-H(b)=H(p_{1}),$
and $g(p_{a})$ continuously decreases from $p_{a}=1-b$ to $p_{a}=1$, we can always find a solution $p_{a}\geq 1-b\geq\frac{1}{2}$ satisfying $g(p_{a})=C$ as long as $H(p_{1})\leq C\leq H(S)$. Hence the value assignment of (11) is valid.
Then we can see that the pair of $(p_{a},p_{b})$ defined in (11) and (10) makes the perception constraint (3a) always inactive, since for $p_{b}=\frac{(1-b)(1-p_{a})}{b}$, we have
$d_{TV}(p_{X},p_{\hat{X}})=|-bp_{b}+(1-b)(1-p_{a})|=0<P.$
At the same time, $H(S|\hat{X})=f(p_{a},\frac{(1-b)(1-p_{a})}{b})=g(p_{a})=C$ for the choice of $p_{a}=g^{-1}(C)$. By Mrs. Gerber’s Lemma[16], when $H(S|\hat{X})\leq C$, we have $H(X|\hat{X})\leq H(C_{1})$, with $C_{1}$ defined as $C_{1}=\frac{H^{-1}(C)-p_{1}}{1-2p_{1}}$. The equality $H(X|\hat{X})=H(C_{1})$ holds when $H(S|\hat{X})=C$. Hence, such choice of $(p_{a},p_{b})$ achieves the lower bound of $I(X;\hat{X})=H(X)-H(C_{1})$ with the perception constraint always inactive.
Appendix D Proof of Theorem 4
Consider the RPC problem as below,
$R(P,C)=$ $\min_{p_{\hat{X}|X}}I(X;\hat{X})$
s.t. $d_{\text{KL}}(p_{X},p_{\hat{X}})\leq P,$
$H(S|\hat{X})\leq C,$
where $(X,S)$ are jointly Gaussian variable with covariance $\text{Cov}(X,S)=\theta_{1}$. Similar to [9, 17], the optimal reconstruction $\hat{X}$ should also be jointly Gaussian with $X$, thus the parameters to be optimized are expectation $\mu_{\hat{x}}$, variance $\sigma_{\hat{x}}^{2}$, and covariance $\text{Cov}(X,\hat{X})=\theta_{2}$.
First we compute the KL-divergence of two Gaussian varibales $X\sim\mathcal{N}(\mu_{x},\sigma^{2}_{x})$ and $\hat{X}\sim\mathcal{N}(\mu_{\hat{x}},\sigma^{2}_{\hat{x}})$. By definition,
$d_{\text{KL}}(p_{X},p_{\hat{X}})$ $=\int p_{X}(x)\log\frac{p_{X}(x)}{p_{\hat{X}}(x)}dx$
[…]
$=\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})+\frac{1}{\sigma_{\hat{x}}^{2}}(\sigma^{2}_{x}+\mu^{2}_{x})+\frac{\mu_{\hat{x}}}{\sigma^{2}_{\hat{x}}}\mu_{x}+\frac{\mu^{2}_{\hat{x}}}{2\sigma^{2}_{\hat{x}}}$
$=\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}}{2\sigma_{\hat{x}}^{2}}.$
Thus,
$d_{\text{KL}}(p_{X},p_{\hat{X}})$ $=\!-\frac{1}{2}\!\log(2\pi e\sigma_{x}^{2})\!+\!\frac{1}{2}\log(2\pi\sigma^{2}_{\hat{x}})\!+\!\frac{(\mu_{x}\!-\!\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}\!+\!\frac{\sigma_{x}^{2}}{2\sigma_{\hat{x}}^{2}}$
$=\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}},$
where we take the natural base of logarithm.
By utilizing the formula of differential entropy and mutual information of the (joint) Gaussian variables[15, Chapter 8], as well as the KL divergence of two Gaussian varibales, the RPC problem can be represented by
$R(P,C)=$ $\min_{\mu_{\hat{x}},\sigma_{\hat{x},}\theta_{2}}-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ (13)
s.t. $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}\leq P,$ (13a)
$-\frac{1}{2}\log(1-\frac{\theta_{1}^{2}}{\sigma_{s}^{2}\sigma_{x}^{4}}\cdot\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}})\geq h(S)-C.$ (13b)
First, we can observe that we can set $\mu_{\hat{x}}=\mu_{x}$ without any effect of the objective function and the constraint (13b), while reducing the perception constraint (13a).
Similar to the RDC case, by (13b) we have
$\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}\geq\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}),$
which gives us a lower bound of the objective function $I(X;\hat{X})$:
$-\frac{1}{2}\log(1\!-\!\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ $\geq\!-\frac{1}{2}\log\Big{(}1\!-\!\frac{1}{\sigma_{x}^{2}}\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1\!-\!e^{-2h(S)+2C})\Big{)}$
$=-\frac{1}{2}\log\Big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\Big{)},$
since $-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})$ is an increasing function of $\frac{\theta_{2}^{2}}{\sigma^{2}_{\hat{x}}}$. The lower bound is achieved when $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$.
Now we show that the lower bound could be achieved with the constraint (13a) always inactive.
Consider the solution
$\sigma_{\hat{x}}=\sigma_{x},$ (14)
$\theta_{2}^{2}=\frac{\sigma_{s}^{2}\sigma_{x}^{6}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C}).$ (15)
Since $\frac{\theta_{2}^{2}}{\sigma_{\hat{x}}^{2}}=\frac{\sigma_{s}^{2}\sigma_{x}^{4}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})$, we have $I(X;\hat{X})=-\frac{1}{2}\log(1-\frac{\theta_{2}^{2}}{\sigma_{x}^{2}\sigma_{\hat{x}}^{2}})=-\frac{1}{2}\log\big{(}1-\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\big{)}$ achieving the lower bound. Meanwhile, we have $\frac{1}{2}\log\frac{\sigma_{\hat{x}}^{2}}{\sigma_{x}^{2}}+\frac{(\mu_{x}-\mu_{\hat{x}})^{2}}{2\sigma_{\hat{x}}^{2}}+\frac{\sigma_{x}^{2}-\sigma_{\hat{x}}^{2}}{2\sigma_{\hat{x}}^{2}}=0<P$ since $\sigma_{\hat{x}}=\sigma_{x}$ and $\mu_{\hat{x}}=\mu_{x}$.
Now we verify that the solution (15) is valid. Since the correlation $\rho=\frac{\theta_{2}}{\sigma_{x}\sigma_{\hat{x}}}$ should have value between -1 and 1, we require that
$\rho^{2}=\frac{\theta_{2}^{2}}{\sigma^{2}_{x}\sigma^{2}_{\hat{x}}}=\frac{\sigma_{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\leq 1,$
which gives us
$C\geq\frac{1}{2}\log\Big{(}1-\frac{\theta_{2}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}\Big{)}+h(S).$
In the RDC Gaussian case, we have proved that that when $C<\frac{1}{2}\log(1-\frac{\theta_{1}}{\sigma_{s}^{2}\sigma_{x}^{2}})+h(S)$, the classification constraint is infeasible. Thus we can safely assume $C\geq\frac{1}{2}\log\Big{(}1-\frac{\theta_{2}^{2}}{\sigma_{s}^{2}\sigma_{x}^{2}}\Big{)}+h(S)$, such that the solution (15) is valid.
Hence, (13a) is always inactive and the optimal value is $I(X;\hat{X})=-\frac{1}{2}\log\big{(}1-\frac{\sigma_{s}^{2}\sigma_{x}^{2}}{\theta_{1}^{2}}(1-e^{-2h(S)+2C})\big{)}$ when $\frac{1}{2}\log(1\!-\!\rho^{2})\!+\!h(S)\leq C\leq h(S)$, solely depending on $C$.